OpenStudy (anonymous):

homework asks to solve by substitution and i can't find any references in my book help? y=3x-16 y=2x-11

OpenStudy (anonymous):

2x-11=3x-16 5=x y=3(5)-16 y=-1 (5,-1)

OpenStudy (anonymous):

thank you very much, it was very helpful the way you explained the procces